Determine algebraically whether the lines are parallel, perpendicular or neither. State the number of solutions. Then draw the graphs of the lines.
x - y= 5
x = 5 - y

Answer by Mathtut(3670)

first write these in point slope form y=mx+b
y=x-5
y=-x+5
:
we know the slope of the 1st is 1 and the slope of the 2nd is -1
since these are negative reciprocals of each other we know the lines are perpendicular
since they are perpendiular to each other there is only one solution and that is where they intersect. (5,0)
I cannot get my graph to work on this so you may have to play around with that.
2 points on the 1st equation are(0,-5) and (5,0) plot and draw a connecting line
2 points on the 2nd equation are (0,5) and (5,0) plot and draw a connecting line